# The Efficient Tail Hypothesis: An Extreme Value Perspective on Market Efficiency

### v1

@@ 3.3 Directional tail dependence measure
The DTD measure can be used to describe the dependence from target $Y$ on the explanatory variable $X$, particularly when focusing on extremal situations. The case $\lambda(X,Y)=0$ implies that the dependence structure for $(X^{+},Y^{+},Y^{-})^{\top}$ is directionally symmetric. We design this measure to satisfy certain desirable properties; see, e.g., Embrechts et al., (2002). Specifically, the DTD measure satisfies the following two properties: first, $\lambda(X,Y)\in[-1,1]$, where $\lambda(X,Y)=-1$ and $\lambda(X,Y)=1$ are achieved when ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1}$ and ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1}$, respectively, which indicate perfect asymptotic dependence between the upper tail of $X$ and, respectively, the lower and upper tails of $Y$; second, $\lambda(X,Y)$ is an odd function of $Y$, that is, $\lambda(X,-Y)=-\lambda(X,Y)$ for all $Y$ (this property does not hold for $X$). Thus, $\lambda(X,Y)$ and $\lambda(-X,Y)$ describe the asymmetry in the influence on $Y$ from the upper tail and lower tail, respectively, of $X$. See the Supplementary Material, Section A.4 for the proof. In summary, the directional tail dependence measure is a flexible description of directional asymmetry in the tails of $(X,Y)^{\top}$, and it is directional, and so has potential usage in causal-effect analysis.
@@
We can view the first form of $\lambda$ in Equation (6) as an extension of the EDM for the vector $(X^{+},Y)^{\top}$, which is regularly-varying on the right half space of $\mathbb{R}^{2}$. The second form of $\lambda$ is the difference between the quadrant-specific measures $\sigma(X^{+},Y^{+})$ and $\sigma(X^{+},Y^{-})$. These two forms are equivalent, but permit two different estimators; for the proof of Proposition 3.4, see the Supplementary Material, Section A.5.
@@ 4.1 Simulation design
$U=U^{\prime},\quad V=\begin{cases}V^{\prime}&\text{, with probability }\phi,\\ 1-V^{\prime}&\text{, otherwise,}\end{cases}$ (9)
@@ 4.2 Results
For each setting, we test the null $\mathcal{H}_{0}:\lambda(X,Y)=0$ for $K=1,000$ repetitions. For each test, we simulate independent samples $\{(x_{i},y_{i})^{\top}\}_{i=1}^{n}$ with sample size $n=10,000$ and calculate the test statistic with threshold $r_{0}$ taken to be the empirical $0.99$-quantile of $\{r_{i}=||(x_{i},y_{i})^{\top}||_{2}\}_{i=1}^{n}$ (and similarly for $r_{0}^{+}$ and $r_{0}^{-}$). We compare the permutation test (with $P=1,000$) for both estimators $\widehat{\lambda}^{1}_{n}$ and $\widehat{\lambda}^{2}_{n}$ in Definition 3.5, with the $t$-test based on the asymptotic normality of $\widehat{\lambda}^{1}_{n}$. In the end we have, for each of the three test types, $K$ test statistics $|\widehat{\lambda}_{n}^{(k)}|$ and their corresponding p-value, $p^{(k)}$, for $k=1,\dots,K$. Figure 3 visualises the relationship between the test statistics and their p-values.
@@
Under the null hypothesis, $p^{(k)}$ should be roughly equal to $1-{\rm rank}(|\widehat{\lambda}^{(k)}_{n})|/K$. In Figure 3, left column-block, we see that all points are clustered around the second diagonal under symmetry, which suggests that all of our estimators perform well and our tests are well calibrated under the null. From the right column-block of Figure 3, which displays the results when data are simulated under the alternative hypothesis (with $\phi=0.7$), all points are below the diagonal line. Points closer to the $x$-axis indicate tests with larger power of rejection. As expected, the statistical power for all three testing methods become stronger with increasing tail asymmetry (descending rows).
@@
We also investigate the effect of the sample size $n$ and lower exceedance thresholds (see Appendix C); the results are similar, and so are not discussed further here. In our application in Section 5, we use the permutation-based test with estimator $\widehat{\lambda}^{2}_{n}$.
@@ 5 Application
5.1 Overview
We apply the framework developed in this paper to analyze the market-wide extremal dependence of China’s futures market. The data are represented as a multivariate time series of percentage price change of futures of commodities with high liquidity. The data are sampled at 30-second intervals and the detailed procedure of data generation and access are given in Appendix B. We consider a subset of the dataset containing 55 assets which covers a wide range of commodities, including agricultural products, base metals, and ferrous metals. We select a one-year period of data from 2022-08-01 to 2023-07-31, which contains $T=\;$$110,985$ observations. To satisfy balanced regular variation (see Definition 3.2), we marginally transform the time series to Symmetric Pareto margins, see (3), using the empirical rank-based transformation given in Equation (10) below:
$X^{\prime}=\begin{cases}(1-\tilde{F}_{X}(X))^{-\frac{1}{2}}-\sqrt{2}&\text{, if }X\geq\tilde{F}_{X}^{-1}(0.5),\\ -(1-\tilde{F}_{-X}(-X))^{-\frac{1}{2}}+\sqrt{2}&\text{, if }X<\tilde{F}_{X}^{-1}(0.5),\\ \end{cases}$ (10)
where $\tilde{F}_{X}(\cdot)$ denotes the empirical distribution function of random variable $X$. We denote these transformed data by $\bm{Z}_{t}=(Z_{1,t},\dots,Z_{55,t})^{\top}$ for $t=1,\dots,T$.
In Section 5.2, we explore the market-wide extremal dependence structure by visualizing dependence on the extremal ball. In Section 5.3, we test the ETH on the market and analyze the results. In Section 5.4, we capitalize on our testing results to construct an artificial dynamic portfolio consisting of tail-inefficient assets and backtest an investment strategy on out-of-sample data.
@@ 5.2 Visualization of market-wide extremal dependence
We are specifically interested in one-step forward extremal dependence. Figure 4 visualises asymmetry in the market-wide extremal dependence for all pairs $(-Z_{i,t},Z_{j,t+1})^{\top}$ and $(Z_{i,t},Z_{j,t+1})^{\top}$ where $i,j=1,\dots,55$. The former group corresponds to those where the driving asset is depreciating, whilst the latter corresponds to pairs where the driving asset is appreciating.
We observe that most of the points concentrate on the tail efficient line, covering a wide range of strengths of extremal dependence. This suggests that most one-step forward extremal dependencies are directionally tail symmetric. A similar visualization is conducted for the contemporaneous extremal dependence structure which reveals the pattern of concurrence of extremal events across assets; we defer this to Appendix D for brevity.
Figure 4: Extremal ball for one-step forward extremal dependencies in China’s futures market. Significant tail-inefficient pairs are circled (based on the permutation test with $\hat{\lambda}^{2}_{n}$ and significance level of 0.01). Red denotes pairs $(Z_{i,t},Z_{j,t+1})^{\top}$ and blue denotes $(-Z_{i,t},Z_{j,t+1})^{\top}$. Arrows point from assets $Z_{i,t}$ (with $\pm$ indicating the direction of $Z_{i,t}$) to assets $Z_{j,t+1}$ for the top 10 most significant tail-inefficient pairs with the largest $|\widehat{\lambda}^{2}_{n}(X_{i},Y_{j})|$ (detailed information is in Table 1).
@@ 5.3 Testing the Efficient Tail Hypothesis for China’s futures market
Although the extremal dependence structure mainly concentrates on the tail efficient line of the extremal ball, one can still notice some deviations. To verify the significance of those deviations, we test the ETH posited in Section 3.4, excluding the last two months (2023-06-01 to 2023-07-31) for an out-of-sample analysis in Section 5.4. We let $\bm{X}=(Z_{1,t},\dots,Z_{55,t},-Z_{1,t},\dots,-Z_{55,t})^{\top}$ be the explanatory vector and $\bm{Y}=(Z_{1,t+1},\dots,Z_{55,t+1})^{\top}$ be the target vector. There are a total of $M=$ $6,050$ pairs $(X_{i},Y_{j})^{\top}$ for $i=1,\dots,110$ and $j=1,\dots,55$. For each pair, we estimate $\lambda(X_{i},Y_{j})$ and its p-value with $P=1000$ permutations using the second proposed estimator $\hat{\lambda}^{2}_{n}$ (see Section 4.2). The quantile for thresholds $r_{0}^{+},r_{0}^{-}$ are derived using the empirical $0.99$-quantiles. The significance level is set to $\alpha^{*}=$ 0.01 (results under $\alpha^{*}=0.05$ are similar) and we use the Benjamini and Hochberg, (1995) multiple testing correction. The p-values for each pair are sorted as $p_{(1)}<\dots<p_{(M)}$. The Benjamini–Hochberg rejection threshold is $p_{(H)}$ with order $H=\max\{i:p_{(i)}<l_{i}\}$, where $l_{i}=\frac{i\alpha^{*}}{C_{M}M}$ for $C_{M}=\sum_{i=1}^{M}i^{-1}$, which accounts for dependence between the tests.
@@
We find that 126 pairs, $(X_{i},Y_{j})^{\top}$, have significant asymmetry, i.e., $\lambda(X_{i},Y_{j})\neq 0$. Therefore, we reject the Efficient Tail Hypothesis for China’s futures market at the significance level of 0.01. These pairs are marked in Figure 4. Here, we focus on analyzing the significant pairs with the top 10 largest test statistic, $|\widehat{\lambda}^{2}_{n}(X_{i},Y_{j})|$; see Table 1. The largest value corresponds to a measure of asymmetry between depreciation (negative) values of red dates (represented by code CJ) and its one-step forward value, with the sign of $\lambda(X_{i},Y_{j})$ being positive. This suggests that extreme depreciation in red dates futures tends to occur with an extreme appreciation in the next period.
Table 1: Asset pairs (and their acronyms) with the top 10 largest asymmetry test statistics.
@@
Explanatory asset $X_{i}$ Target asset $Y_{j}$ Sign of $X_{i}$ $\widehat{\lambda}^{2}_{n}(X_{i},Y_{j})$
red dates (CJ) red dates (CJ) negative 0.073
coking coal (jm) coke (j) negative -0.065
iron ore (i) coking coal (jm) negative -0.063
soybean oil (y) vegetable oil (OI) positive 0.057
low sulfur fuel (lu) cotton yarn (CY) positive 0.056
coke (j) coking coal (jm) negative -0.055
coking coal (jm) iron ore (i) negative -0.055
styrene (eb) methanol (MA) negative -0.055
polypropylene (pp) methanol (MA) negative -0.054
iron ore (i) soda ash (SA) negative -0.053
@@
We find that eight out of those ten aforementioned pairs are driven by negative extreme events. This suggests that the extremal depreciation of assets are more influential than the extremal appreciation of assets. Similar conclusions are found by Bouaddi and Moutanabbir, (2022), who find that negative extreme movements are more influential than positive extreme movements in the U.S. stock market.
@@
By analyzing the sign of the explanatory variable $X_{i}$ and the sign of $\lambda(X_{i},Y_{j})$, we can classify the significant pairs into two types: momentum effects, where ${\rm Sgn}(X_{i})={\rm Sgn}(\lambda(X_{i},Y_{j}))$, and contrarian effects, otherwise. Momentum (contrarian) effects are pervasive phenomena in financial markets, reflecting a tendency for price changes to follow the same (different) directional price movements; see e.g., Asness et al., (2013) for more details. Nine out of the top ten significant pairs are driven by momentum effects. The phenomenon of momentum in extremes may suggest that the market is under-active in extreme situations and indicates the potential contagion path of extreme risk when the market undergoes stress. The phenomenon of contrarian in extremes may suggest that the market is over-active in extreme situations and tends to recover in the following periods.
@@
We also analyze the in-product (i.e., the same asset for $X_{i}$ and $Y_{j}$ ) and cross-product (i.e., different assets) pairs. Table 1 shows that the majority of the significant pairs are cross-product pairs (9 out of 10) and most pairs bridge assets in similar sectors with one exception (the pair for low sulfur fuel and cotton yarn).
@@
Statistics on all 126 significant pairs have similar patterns: 68 pairs are driven by negative extreme events and 58 pairs are driven by positive extreme events; 102 pairs are driven by momentum effects and 24 pairs are driven by contrarian effects. There are only three in-product significant pairs: red dates (driven by negative extreme events and is contrarian), coke (driven by negative extreme events and is momentum), and vegetable oil (driven by positive extreme events and is contrarian). The rest of the 123 pairs are cross-product pairs. These results suggest that negative extreme movements are more influential than positive extreme movements, and that momentum effects in extremes tend to happen in cross-product pairs while contrarian effects tend to happen in in-product pairs.

### v2

@@ 3.4 Directional tail dependence measure
The DTD measure can be used to describe the dependence from target $Y$ on the explanatory variable $X$, particularly when focusing on extremal situations. The case $\lambda(X,Y)=0$ implies that the dependence structure for $(X^{+},Y^{+},Y^{-})^{\top}$ is directionally symmetric. We design this measure to satisfy certain desirable properties; see, e.g., Embrechts et al., (2002). Specifically, the DTD measure satisfies the following two properties as stated in Proposition 3.4 (see the Supplementary Material, Section A.4 for the proof).
Proposition 3.4 (Properties of directional tail dependence $\lambda$).
The directional tail dependence $\lambda(X,Y)$ satisfies ${\lambda(X,Y)\in[-1,1]}$, where ${\lambda(X,Y)=-1}$ and ${\lambda(X,Y)=1}$ are achieved when ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1}$ and ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1}$, respectively. Furthermore, $\lambda(X,Y)$ is an odd function of $Y$, i.e., $\lambda(X,-Y)=-\lambda(X,Y)$ for all $Y$ (this property does not hold for $X$).
The boundary cases ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1}$ and ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1}$ correspond to perfect asymptotic dependence between the upper tail of $X$ and, respectively, the lower and upper tails of $Y$. Moreover, $\lambda(X,Y)$ is an odd function of $Y$ only. Thus, $\lambda(X,Y)$ and $\lambda(-X,Y)$ describe the asymmetry in the influence on $Y$ from the upper tail and lower tail, respectively, of $X$. In summary, the directional tail dependence measure is a flexible description of directional asymmetry in the tails of $(X,Y)^{\top}$, and it is directional, and so has potential usage in causal-effect analyses.
@@
We can view the first form of $\lambda$ in Equation (7) as an extension of the EDM for the vector $(X^{+},Y)^{\top}$, which is regularly varying on the right half space of $\mathbb{R}^{2}$. The second form of $\lambda$ is the difference between the quadrant-specific measures $\sigma(X^{+},Y^{+})$ and $\sigma(X^{+},Y^{-})$. These two forms are equivalent, but permit two different estimators; for the proof of Proposition 3.5, see the Supplementary Material, Section A.5.
@@ 4.1 Simulation design
$U=U^{\prime},\quad V=\begin{cases}V^{\prime}&\text{with probability }\phi,\\ 1-V^{\prime}&\text{otherwise,}\end{cases}$ (10)
@@ 4.2 Results
For each setting, we test the null $\mathcal{H}_{0}:\lambda(X,Y)=0$ for $K=1,000$ repetitions. For each test, we simulate independent samples $\{(x_{i},y_{i})^{\top}\}_{i=1}^{n}$ with sample size $n=10,000$ and calculate the test statistic with threshold $r_{0}$ taken to be the empirical $0.99$-quantile of $\{r_{i}=||(x_{i},y_{i})^{\top}||_{2}\}_{i=1}^{n}$ (and similarly for $r_{0}^{+}$ and $r_{0}^{-}$). We compare the permutation test (with $P=1,000$) for both estimators $\widehat{\lambda}^{1}_{n}$ and $\widehat{\lambda}^{2}_{n}$ in Definition 3.5, with the $t$-test based on the asymptotic normality of $\widehat{\lambda}^{1}_{n}$. In the end, for each of the three test types, we have $K$ independent pairs of test statistics $|\widehat{\lambda}_{n}^{(k)}|$ and their corresponding p-value, $p^{(k)}$, for $k=1,\dots,K$. Figure 3 visualises the relationship between the test statistics and their p-values.
@@
Under the null hypothesis, $p^{(k)}$ should be roughly equal to $1-{\rm rank}(|\widehat{\lambda}^{(k)}_{n})|/K$. In Figure 3, left column-block, we see that all points are clustered around the second diagonal under symmetry, which means that the rejection percentages for each significance level are approximately equal to the corresponding nominal value. From the right column-block of Figure 3, which displays the results when data are simulated under the alternative hypothesis (with $\phi=0.7$), all points are below the diagonal line. Points closer to the $x$-axis indicate tests with larger power of rejection. As expected, the statistical power for all three testing methods become stronger with increasing tail asymmetry (descending rows).
@@
We also investigate the effect of the sample size $n$ and lower exceedance thresholds (see the Supplementary Material, Section C); the results are similar, and so are not discussed further here. In our application in Section 5, we use the permutation-based test with estimator $\widehat{\lambda}^{2}_{n}$.
@@ 5 Application
5.1 Overview and basic data analysis
We apply the framework developed in this paper to analyze the market-wide extremal dependence of China’s futures market. The data are represented as a multivariate time series of percentage price changes of futures for commodities with high liquidity. The return is calculated at 30-second intervals and the detailed procedure of data generation and access are given in the Supplementary Material, Section B. We consider a subset of the dataset containing 55 assets, which covers a wide range of commodities, including agricultural products, base metals, and ferrous metals. We select a one-year period of data from 2022-08-01 to 2023-07-31, which contains $T=\;$$110,985$ observations.
For each of the $55$ assets, we study bulk and tail temporal dependence by estimating their autocorrelation and extremogram (Davis et al.,, 2012) functions, respectively. The estimated (extremal) dependence decays rapidly, dropping below values of $0.05$ and $0.1$ (for the auto-correlation and extremogram, respectively) after $10$ time lags. Therefore, to remove temporal dependence from the data, we subsample every tenth observation. A detailed description of this dependence analysis is deferred to the Supplementary Material, Section D.1. After sub-sampling, we transform the data to be balanced regularly varying. In the Supplementary Material, Section D.2, we investigate the marginal tail behaviour of the data and find reasonable evidence to suggest that the upper and lower tails of each asset are regularly varying; we thus employ the first marginal transformation as described in Section 3.2. Unreported experiments found this choice to have negligible effect on the final results. We denote the transformed data by $\bm{Z}_{t}=(Z_{1,t},\dots,Z_{55,t})^{\top}$ for $t=1,\dots,T$.
The rest of this section is organized as follows: In Section 5.2, we visualize extremal dependence in the market using the extremal ball. In Section 5.3, we test the ETH for China’s futures market and interpret our findings. In Section 5.4, we capitalize on our test results to construct an artificial dynamic portfolio consisting of tail-inefficient assets and backtest an investment strategy on out-of-sample data. We also study the uncertainty arising from the marginal transformation via a bootstrap experiment, which is deferred to Supplementary Material, Section D.4.
@@ 5.2 Visualization of market-wide extremal dependence
We are specifically interested in one-step forward extremal dependence, i.e., the extremal dependence in the bivariate vectors $(-Z_{i,t},Z_{j,t+1})^{\top}$ and $(Z_{i,t},Z_{j,t+1})^{\top}$ where $i,j=1,\dots,55$. Figure 4 visualises directional tail asymmetry in the market-wide extremal dependence for all pairs $(-Z_{i,t},Z_{j,t+1})^{\top}$ and $(Z_{i,t},Z_{j,t+1})^{\top}$ where $i,j=1,\dots,55$. The former group corresponds to those where the driving event is extremal loss, whilst the latter corresponds to pairs where the driving asset is the extremal gain.
We observe that most of the points concentrate on the tail efficient line, covering a wide range of strengths of extremal dependence. This suggests that most one-step forward extremal dependencies are directionally tail symmetric. Although some extreme events appear to be closely linked to subsequent large movements in asset prices, this connection does not necessarily yield a clear directional signal. At best, it serves as an indicator of heightened volatility, implying that the market remains relatively efficient even under extreme conditions. A similar visualization is conducted for the contemporaneous extremal dependence structure which reveals the pattern of concurrence of extremal events across assets; we defer this to Supplementary Material, Section D.3 for brevity.
Figure 4: Extremal ball for one-step forward extremal dependencies in China’s futures market. Significant tail-inefficient pairs with p-values smaller than 0.01 are circled (based on the permutation test with $\hat{\lambda}^{2}_{n}$). Red points denote pairs $(Z_{i,t},Z_{j,t+1})^{\top}$ and blue points denote $(-Z_{i,t},Z_{j,t+1})^{\top}$.
@@ 5.3 Testing the Efficient Tail Hypothesis for China’s futures market
Although the extremal dependence structure mainly concentrates on the tail efficient line of the extremal ball, one can still notice some deviations. To verify the significance of those deviations, we test the ETH posited in Section 3.5, excluding the last two months (2023-06-01 to 2023-07-31) for an out-of-sample analysis in following Section 5.4. We let $\bm{X}=(Z_{1,t},\dots,Z_{55,t},-Z_{1,t},\dots,-Z_{55,t})^{\top}$ be the explanatory vector and asset return after one period $\bm{Y}=(Z_{1,t+1},\dots,Z_{55,t+1})^{\top}$ be the target vector. There are a total of $M=$ $6,050$ pairs $(X_{i},Y_{j})^{\top}$ for $i=1,\dots,110$ and $j=1,\dots,55$. For each pair, we estimate $\lambda(X_{i},Y_{j})$ and its p-value with $P=10,000$ permutations using the second proposed estimator $\hat{\lambda}^{2}_{n}$ (see Section 4.2). The quantile for thresholds $r_{0}^{+},r_{0}^{-}$ are derived using the empirical 0.99-quantiles. The significance level is set to $\alpha^{*}=$0.01 and we use the Benjamini and Hochberg, (1995) multiple testing correction. The p-values for each pair are sorted as $p_{(1)}<\dots<p_{(M)}$. The Benjamini–Hochberg rejection threshold is $p_{(H)}$ with order $H=\max\{i:p_{(i)}<l_{i}\}$, where $l_{i}=\frac{i\alpha^{*}}{C_{M}M}$ for $C_{M}=\sum_{i=1}^{M}i^{-1}$, which accounts for dependence between the tests. The Efficient Tail Hypothesis (ETH) for China’s futures market, that is, that the pairwise ETH is rejected for at least one pair of assets (see Definition 3.6), is rejected under the significance level of 0.01. This result is consistent across the 100 bootstrap replicates, where the market-level ETH is rejected for all replicates.
@@
By studying pairs with p-values smaller than 0.01, we find that 60 pairs are tail-inefficient. These pairs are marked in Figure 4. Here, we focus on analyzing the significant pairs with the top 10 largest test statistic, $|\widehat{\lambda}^{2}_{n}(X_{i},Y_{j})|$; see Table 1. Table 1 also shows the bootstrap mean and standard deviation of $\widehat{\lambda}^{2}_{n}(X_{i},Y_{j})$ across $100$ bootstrap replicates. The largest value corresponds to a measure of directional asymmetric dependence between the eggs’ futures and itself in the next time step, with a p-value of $0.0051$. This suggests that extreme depreciation in futures of eggs tends to occur with an extreme appreciation in the next period.
Table 1: Asset pairs (and their acronyms) with the top 10 largest asymmetry test statistics for pairs with p-values smaller than 0.01. The bootstrap mean and standard deviation of $\hat{\lambda}^{2}_{n}$ are calculated using $100$ bootstrap replicates. Each of the estimated $\hat{\lambda}^{2}_{n}$ for the observed data are within one standard deviation of the bootstrap mean, and omitted.
@@
Explanatory asset $X_{i}$ Target asset $Y_{j}$ Sign of $X_{i}$ p-values Boot. mean $\widehat{\lambda}^{2}_{n}$ (std)
eggs (jd) eggs (jd) negative 0.0051 0.145 (0.054)
zinc (zn) fuel oil (fu) negative 0.0012 -0.130 (0.044)
live hog (lh) pulp (sp) negative 0.0001 -0.127 (0.036)
Polypropylene (pp) Polyvinyl chloride (v) negative 0.0018 -0.120 (0.036)
corn starch (cs) aluminum (al) positive 0.0002 -0.113 (0.052)
pulp (sp) red dates (CJ) negative 0.0006 0.111 (0.035)
Ethylene Glycol (eg) Soda Ash (SA) negative 0.0004 -0.108 (0.036)
corn starch (cs) lead (pb) negative 0.0011 -0.105 (0.055)
eggs (jd) lead (pb) positive 0.0018 -0.092 (0.063)
natural rubber (ru) natural rubber (ru) positive 0.0063 -0.074 (0.064)
@@
We find that seven out of the ten aforementioned pairs are driven by negative extreme events. This suggests that the extremal depreciation of assets are more influential than the extremal appreciation of assets. Similar conclusions are found by Bouaddi and Moutanabbir, (2022), who show that negative extreme movements are more influential than positive extreme movements in the U.S. stock market.
@@
By analyzing the sign of the explanatory variable $X_{i}$ and the sign of $\lambda(X_{i},Y_{j})$, we can classify the significant pairs into two types: momentum effects, where ${\rm Sgn}(X_{i})={\rm Sgn}(\lambda(X_{i},Y_{j}))$, and contrarian effects, otherwise. Momentum (contrarian) effects are well-known phenomena in financial markets, reflecting a tendency for price changes to follow the same (different) directional price movements; see e.g., Asness et al., (2013) for more details. The phenomenon of momentum in extremes may suggest that the market is under-active in extreme situations and indicates the potential contagion path of extreme risk when the market undergoes stress. The phenomenon of contrarian in extremes may suggest that the market is over-active during extreme events and tends to recover in the following periods.
@@
We also analyze the in-product (i.e., the same asset for $X_{i}$ and $Y_{j}$ ) and cross-product (i.e., different assets) pairs. Table 1 shows that the majority of the significant pairs are cross-product pairs (8 out of 10).
@@
Statistics on all 60 significant pairs exhibit similar patterns: 41 pairs are driven by negative extreme events and 19 pairs by positive extreme events; 41 pairs are driven by momentum effects and 19 by contrarian effects. There are only three in-product significant pairs: egg futures (driven by an extremal loss), natural rubber (driven by an extremal gain), and stainless steel (driven by an extremal gain). These three pairs are all contrarian in extremes, while the remaining 57 pairs are cross-product pairs. Within the cross-product pairs, 41 pairs are driven by momentum effects.
Overall, the findings suggest that negative extreme movements are more influential than positive ones, and that momentum effects in extremes tend to appear in cross-product pairs whereas contrarian effects tend to appear in in-product pairs.
Such statistics show that, in the context of extremes of China’s futures market, an asset’s own price often exhibits an initial overreaction, followed by a subsequent correction, which may form the basis of contrarian effects in extremes for tail-inefficient in-product pairs. As for the momentum effects in extremes for the cross-product pairs, as mentioned in Hou, (2007), the speed at which the price of an asset reacts to new information varies among products. The “slow information diffusion hypothesis”, as described by Hou, (2007), may explain our observed momentum effects in the cross-product pairs. One plausible explanation is that certain assets swiftly incorporate relevant news, leading to initial price adjustments in explanatory assets, while others reflect the same information only after a delay. However, it remains an open question as to whether the observed cross-product linkages in extremes reflect truly causal relationships.